OpenStudy (anonymous):

Suppose the vertex angle in an equiangular triangle is represented by 2x + 32. What is the value of x?

OpenStudy (anonymous):

3(2x+32) = 180 12X + 96= 180 12x = 180-96 12x = 84 x = 7

OpenStudy (anonymous):

That's incorrect. The only option for this question is: A- 60 B- 28 C- 14 D- 32

OpenStudy (anonymous):

hmm x is 7 , wait @TuringTest @satellite73 @dpaInc